the dimension of a rectangular prism are n,n+7 and n+8.use the formula v=lwh to write a polynomial in standard form for the volume of the prism
Answer by ewatrrr(24785)

 
Hi,
the dimension of a rectangular prism are n,n+7 and n+8.
V(rectangular prism) = lhw = n(n+7)(n+8) = n(n^2 + 15n + 56) = n^3 +15n^2 + 56n